# Two-point correlators in de Sitter-prepared states with bra-ket wormholes

### v1

@@ Abstract
Motivated by the finiteness of de Sitter (dS) horizon entropy, we study how “bra-ket wormholes” modify correlation functions in gravitationally prepared states. Euclidean wormhole saddles in gravitational path integrals can generate non-factorizing contributions to correlation functions, as in replica-wormhole explanation of the Page curve and bra-ket-wormhole restoration of strong subadditivity. By defining ‘time’ variables and computing observables in a flat region attached to the dS boundary, we evaluate bra–ket wormhole contributions to scalar two-point functions and find a late-time transition in the dominant saddle, accompanied by ramp-and-plateau behavior of correlations and a characteristic timescale comparable to the fast scrambling. Our results are built upon (i) inflationary horizon exit and re-entry, (ii) enhancement of correlation at small comoving momentum $k$ by wormhole contribution, and (iii) a competition between mode counting and topological suppression that drives a transition to wormhole dominance. Although results are qualitatively consistent, one needs to address wormhole stabilization to clearly interpret in terms of the entropy context.
@@ 1 Introduction
The gravitational path integral prepares a quantum state on the boundary of quasi-dS—reheating surface—by integrating over all possible past Euclidean geometries and field configurations given boundary conditions Hartle:1983ai. Although past time slices in path integral cannot be interpreted as real dynamics of inflation and observables inside dS are uncertain Witten:2001kn; Castro:2012gc, the state can be measured in a post-reheating flat region as discussed. The two observables based on different notions of time allow complementary scrutinizations of the state.
In a finite-entropy system, late-time correlators are not expected to decay indefinitely; rather, they should exhibit residual correlations Maldacena:2001kr and, after appropriate averaging, ramp-and-plateau behavior Saad:2018bqo; Saad:2019lba; Saad:2019pqd. A natural scale for the onset of such deviations is a dS analogue of the Page time: irreducible errors due to the finite number of states $e^{-S_{\mathrm{dS}}}$ ($\propto e^{-1/G}$, hence non-perturbative) can become order one when enhanced by an effective multiplicity of available states of order $e^{S_{\mathrm{dS}}}$ Page:1993wv; Arkani-Hamed:2007ryv. However, horizon is also thought to be a chaotic fast-scrambling system, out of which information can leak far earlier than the Page time Hayden:2007cs; Susskind:2011ap; Maldacena:2015waa. Determining the mechanism that supplies these non-perturbative corrections and how it manifests in correlators of cosmological interest are the central motivations of this work.
A concrete precedent comes from the black hole information problem Hawking:1975vcx; Preskill:1992tc, where replica and bra-ket wormhole saddles modify semiclassical calculations and yield late-time behaviors consistent with unitarity Almheiri:2020cfm; Penington:2019kki. Inspired by this mechanism, we study whether analogous bra-ket wormholes Chen:2020tes; Fumagalli:2024msi contribute to correlators in dS-prepared states, in ways consistent with the finite-entropy constraints.
@@ 2 Gravitationally prepared state at the end of inflation
Figure 1: dS prepares a quantum state (of the scalar field) on its spatial boundary $\Sigma$, reheating surface, including effects from higher-topology wormholes. The state is then observable in the flat space afterwards without gravity. The observation utilizes two notions of time: $k_{\rm max}$ and $k_{\rm min}$.
@@ 2.2 Observation of dS prepared state, notion of time
Post-inflationary observation. Our situation in Fig. 1 is interpreted as inflation prepares a quantum state of the universe and scalar field (representing inflaton or the energy density fluctuation) on the reheating surface $\Sigma$, which is subsequently evolved into the post-inflationary FRW space. This state is observable by post-inflationary observers, e.g. by CMB and large scale structures.
@@
From the so called CMB observer’s point of view, observed properties of dS state change with time, as successively smaller Fourier modes of quantum fields (larger wavelength) re-enter horizon. The measured entropy also grows, and eventually becomes conflict with the finite dS entropy. Presumably, wormhole effects would cure this conflict, just like in the black hole case. In Sec. 4.3, we will show that wormhole effects encoded in various Fourier modes of dS state indeed make CMB observation consistent with unitarity and complementarity.
@@ 3.1 Equations of motion and boundary conditions in the global coordinate
$\Phi_{k}(t)\,\simeq\,\frac{\phi(k)}{u(t_{\epsilon})}\,\left(\phi(k)\,a(t)^{-\Delta_{-}}+\mathcal{O}(k)\,a(t)^{-\Delta_{+}}\right),\qquad\Delta_{\pm}\,\equiv\,\frac{1}{2}\pm\sqrt{\frac{1}{4}-m^{2}},$ (21)
@@ 3.3 Bra-ket wormhole by tracing out unobservable universe
One may wonder if our mechanism of forming an effective bra-ket wormhole by joining bra-bra and ket-ket wormholes can be extended by joining even more chains of bra-bra and ket-ket wormholes. They are indeed legible contributions. But, as detailed in Appendix D, adding $n$ such chains increases the action in proportion to $\propto n$ so that they are exponentially suppressed. Creating longer and more complex wormholes essentially costs more resources. We ignore these contributions.
@@ 3.4 Stabilization of wormholes
In the random state model of quantum gravity Jung:2025RMPS, wormholes indeed turn out to be stabilized; without encountering any divergences, this model successfully reproduces the expected properties of correlation functions in the wormhole dominant phases. Also, in Ref. Fumagalli:2024msi, Wigner distribution of observables also provided some stabilizing force. Accepting these general expectations, we will simply assume a non-zero finite value of $\tau_{0}=0.01$ in our numerical calculations. But true saddle points shall be in relation to $S_{\rm dS}$ as will be discussed, which we leave for future works.
@@
However, this mechanism of stabilization cannot be complete as it does not depend on the gravity action. This will be discussed later in Fig. 10. Thus, we do not take this resolution with extra fermions seriously. Rather, in this study, we simply take the fixed value $\tau_{0}=10^{-2}$.
@@ 4.1 Notions of time: $k_{\rm max}$ and $k_{\rm min}$
In post-inflationary observation of dS prepared state, only those modes of scalar field that became superhorizon by the end of inflation will be classicalized and observable. We introduce upper limit on the observable Fourier modes of scalar field, $k_{\rm max}$, as the largest $k$-mode that existed horizon ($1/H$) by the end of inflation
@@
Observable 2, on the other hand, considers CMB-like two-point correlations, hence only among those modes that have successively re-entered horizon by given time. This introduces another notion of time: $k_{\rm min}$ as the smallest Fourier modes of scalar fields that have re-entered by given time
@@
with $\langle\phi_{k}\phi_{-k}\rangle$ given by Eq. (42). The farther future from reheating, the smaller $k_{\rm min}$ and the more modes have re-entered horizon. By varying $k_{\rm min}$, we really mean to study the time-dependence of CMB observation of the given prepared state by a single observer.
@@ 4.2 Phase transition to wormhole dominance
We denote the critical point by $k_{\rm max}=k_{\rm max}^{c}$, where $Z_{\rm rat}=e^{-S_{\rm dS}}$. This is where phase transition occurs.
@@
In our bra-ket wormhole solution, as discussed in Sec. 3.3, ${\cal A}_{k}^{WH}<{\cal A}_{k}^{HH}$ for small $k<k^{*}$ indeed; large-$k$ modes shall agree between wormhole and HH. Briefly recapping, bra-bra wormhole-induced mixing (or entanglement) between scalar fields in separate universes
$a_{k}\phi_{1}^{2}(k)+2b_{k}\phi_{1}(k)\phi_{2}(k)-a_{k}\phi_{2}^{2}(k)\,\to\,-\Bigl(a_{k}-\frac{b_{k}^{2}}{a_{k}}\Bigr)\phi_{1}^{2}(k)$ (47)
@@
effectively enhances low-$k$ mode correlations $\sim 1/a_{k}^{\rm eff}$ of $\phi_{1}$, by effectively reducing $a_{k}\to a_{k}^{\rm eff}=(a_{k}^{2}-b_{k}^{2})/a_{k}$. For this to work, $b_{k}<a_{k}$, consistently with KSW.
Although we numerically confirmed this low-$k$ enhancement, it is not straightforward to answer why this is compelling. But tracing out the half of dS boundary space is known to induce such an enhancement of low-$k$ correlation in the remaining subsystem Kanno:2014ifa. Interactions between two CFTs (that represent each side of a maximally symmetric black hole or each boundary of AdS) can be responsible for the formation of traversable wormholes Gao:2016bin; Maldacena:2017axo; Maldacena:2018lmt. Thermofield double state is also known to be formed by double trace deformation Cottrell:2018ash. ER=EPR is another related conjecture Maldacena:2013xja; Susskind:2014yaa. All these may support why the enhancement of low-$k$ correlation is compelling from interactions between two universes in our solution.

### v2

@@ Abstract
Motivated by the finiteness of de Sitter (dS) horizon entropy, we study how “bra-ket wormholes” modify correlation functions in gravitationally prepared states. Euclidean wormhole saddles in gravitational path integrals can generate non-factorizing contributions to correlation functions, as in replica-wormhole explanation of the Page curve and bra-ket-wormhole restoration of strong subadditivity. By defining ‘time’ variables and computing observables in a flat region attached to the dS boundary, we evaluate bra–ket wormhole contributions to scalar two-point functions and find late-time transitions in the dominant saddle, accompanied by the ramp-and-plateau behavior of correlations and the characteristic timescale comparable to the fast scrambling. Each observable is consistent with ‘complementarity’, in the sense that wormhole effects are distinguishable only at late respective times. Consistencies are based upon the interplay of (i) inflationary horizon exit and re-entry, (ii) enhancement of correlations at small comoving momentum by wormhole contributions, (iii) a competition between mode counting and topological suppression that drives a transition to wormhole dominance, which naturally yields the fast scrambling timescale, and (iv) irreducible errors by cosmic variance in early CMB-like observations. To clearly interpret in terms of entropy and chaotic nature of dS, one needs a more complete mechanism of wormhole stabilization.
@@ 1 Introduction
The gravitational path integral prepares a quantum state on the boundary of quasi-dS—reheating surface—by integrating over all possible past Euclidean geometries and field configurations given boundary conditions Hartle:1983ai. Although past time slices in the path integral cannot be interpreted as real dynamics of inflation and observables inside the dS are ambiguous Witten:2001kn; Castro:2012gc, the state can be unambiguously measured in a post-reheating flat region, providing a diagnosis of the dS quantum gravity. The two observables based on each notion of time allow complementary scrutinizations of the dS-prepared state.
In a finite-entropy system, late-time correlators are not expected to decay indefinitely; rather, they should exhibit residual correlations Maldacena:2001kr and, after appropriate averaging, ramp-and-plateau behavior Saad:2018bqo; Saad:2019lba; Saad:2019pqd. A natural scale for the onset of such deviations is a dS analogue of the Page time: irreducible errors due to the finite number of states $e^{-S_{\mathrm{dS}}}$ ($\propto e^{-1/G}$, hence non-perturbative) can become order one when enhanced by an effective multiplicity of available states of order $e^{S_{\mathrm{dS}}}$ Page:1993wv; Arkani-Hamed:2007ryv. However, a horizon is also thought to be a chaotic fast-scrambling system, out of which information can leak far earlier than the Page time Hayden:2007cs; Susskind:2011ap; Maldacena:2015waa. Unlike the entanglement entropy, correlation functions can indeed be sensitive to wormhole effects at the scrambling time, e.g., via the scrambling shock-wave protocol Shenker:2013pqa; Gao:2016bin; Maldacena:2017axo; Aalsma:2021kle, although its generalization is unclear. Determining the mechanism that supplies these non-perturbative corrections and how it manifests in correlators of cosmological interest are the central motivations of this work.
A concrete precedent comes from the black hole information problem Hawking:1975vcx; Preskill:1992tc, where replica and bra-ket wormhole saddles modify semiclassical calculations and yield late-time behaviors consistent with unitarity Almheiri:2020cfm; Penington:2019kki. Inspired by this mechanism, we study whether analogous bra-ket wormholes Chen:2020tes; Fumagalli:2024msi contribute to correlators in dS-prepared states, in ways consistent with finite-entropy constraints.
@@ 2 Gravitationally prepared state at the end of inflation
Figure 1: dS prepares a quantum state (of the scalar field) on its spatial boundary $\Sigma$, reheating surface, including effects from higher-topology wormholes. The state is then observable in the flat space afterwards without gravity. Two notions of time, $k_{\rm max}$ and $k_{\rm min}$, define two observables.
@@ 2.2 Observation of dS prepared state, notion of time
Post-inflationary observation. Our situation in Fig. 1 is interpreted as inflation prepares a quantum state of the universe and the scalar field (representing the inflaton or the energy density fluctuation) on the reheating surface $\Sigma$, which is subsequently evolved into the post-inflationary FRW space. This state is observable by post-inflationary observers, e.g. by CMB and large scale structures.
@@
From the so called CMB observer’s point of view, observed properties of dS state change with time, as successively smaller Fourier modes of quantum fields (larger wavelength) re-enter the horizon. The measured entropy also grows, and eventually becomes conflict with the finite dS entropy. Presumably, wormhole effects would cure this conflict, just like in the black hole case. In Sec. 4.4, we will show that wormhole effects encoded in various Fourier modes of dS state indeed make CMB observation at least qualitatively consistent with unitarity and complementarity.
@@ 3.1 Equations of motion and boundary conditions in the global coordinate
$\Phi_{k}(t)\,\simeq\,\frac{\phi(k)}{u(t_{\epsilon})}\,\left(\phi(k)\,a(t)^{-\Delta_{-}}+\mathcal{O}(k)\,a(t)^{-\Delta_{+}}\right),\qquad\Delta_{\pm}\,\equiv\,\frac{1}{2}\pm\nu,\quad\nu\,\equiv\,\sqrt{\frac{1}{4}-m^{2}},$ (21)
@@ 3.3 Bra-ket wormhole by tracing out unobservable universe
One may wonder if our mechanism of forming an effective bra-ket wormhole by joining bra-bra and ket-ket wormholes can be extended by joining even more chains of bra-bra and ket-ket wormholes. They are indeed legible contributions. But, as detailed in Appendix C, adding $n$ such chains increases the action in proportion to $\propto n$ so that they are exponentially suppressed. Creating longer and more complex wormholes essentially costs more resources. We ignore these contributions.
@@ 3.4 Stabilization of wormholes
In the random state model of quantum gravity Jung:2025rip, wormholes indeed turn out to be stabilized; without encountering any divergences, this model successfully reproduces the expected properties of correlation functions in the wormhole dominant phases. Also, in Ref. Fumagalli:2024msi, Wigner distribution of observables also provided some stabilizing force. Accepting these general expectations, we will simply assume a non-zero finite value of $\tau_{0}=0.01$ in our numerical calculations. But true saddle points shall be in relation to $S_{\rm dS}$ as will be discussed, which we leave for future works.
@@
However, this mechanism of stabilization cannot be complete as it does not depend on the gravity action. Thus, we do not take this resolution with extra fermions seriously. Rather, in this study, we simply take the fixed value $\tau_{0}=0.01$.
@@ 4.1 Notions of time: $k_{\rm max}$ and $k_{\rm min}$
In post-inflationary observation of dS prepared state, only those modes of scalar field that became superhorizon by the end of inflation will be classicalized and observable. We introduce upper limit on the observable Fourier modes of scalar field, $k_{\rm max}$, as the largest $k$-mode that existed the horizon ($1/H$) by the end of inflation
@@
Observable 2, on the other hand, considers CMB-like two-point correlations, hence only among those modes that have successively re-entered the horizon by given time. This introduces another notion of time: $k_{\rm min}$ as the smallest Fourier modes of scalar fields that have re-entered by given time
@@
with $\langle\phi_{k}\phi_{-k}\rangle$ given by Eq. (42). The farther future from reheating, the smaller $k_{\rm min}$ and the more modes have re-entered the horizon. By varying $k_{\rm min}$, we really mean to study the time-dependence of CMB observation of the given prepared state by a single observer.
@@ 4.2 Phase transition to wormhole dominance
We denote the critical point by $k_{\rm max}^{c}$, where $Z_{\rm rat}=e^{-S_{\rm dS}}$. This is where phase transition occurs in the full state, or in Observable 1. When we say phase transition, we usually refer to this one. Another phase transition in Observable 2, $k_{\rm min}^{c}$, will also be introduced later, based on the same mechanism while realized slightly differently.
@@
In our bra-ket wormhole solution, the necessary condition ${\cal A}_{k}^{WH}<{\cal A}_{k}^{HH}$ is satisfied for small $k<k^{*}$ as discussed in Sec. 3.3; for large $k>k^{*}$, WH contributions agree with HH. Briefly recapping, bra-bra wormhole-induced mixing (or entanglement) between scalar fields in separate universes
$a_{k}\phi_{1}^{2}(k)+2b_{k}\phi_{1}(k)\phi_{2}(k)-a_{k}\phi_{2}^{2}(k)\quad\to\quad-\Bigl(a_{k}-\frac{b_{k}^{2}}{a_{k}}\Bigr)\phi_{1}^{2}(k)$ (47)
@@
effectively enhances low-$k$ mode correlations $\sim 1/a_{k}^{\rm eff}$ of $\phi_{1}$, by effectively reducing $a_{k}\to a_{k}^{\rm eff}=(a_{k}^{2}-b_{k}^{2})/a_{k}$. For this to work, $b_{k}<a_{k}$, which is consistent with the relaxed KSW.
The resulting enhancement of low-$k$ correlations are numerically confirmed in later sections. This is necessary for the KSW as well as phase transitions into wormhole dominance. However, whether and how these are due to the structure of our bra-ket wormhole (tracing out the second universe) is not clear to us. At least, tracing out the half of the dS boundary space is known to induce such an enhancement of low-$k$ correlations in the remaining subsystem Kanno:2014ifa. Wormholes may induce entangled states, e.g. as in ER=EPR conjecture Maldacena:2013xja; Susskind:2014yaa, or in another way, interactions may induce wormholes Gao:2016bin; Maldacena:2017axo; Maldacena:2018lmt or entangled states Cottrell:2018ash. The scrambling physics that will be discussed in next subsections may also be relevant.
4.3 Parameter space: Scaling and Scrambling
Figure 4: The scaling relation between free parameters of the model, $\{S_{\rm dS},\tau_{0},k_{\rm max}\}$, using $k_{\rm max}^{c}$ as a function of $\tau_{0}$ for two choices of $S_{\rm dS}=76$(solid), 760(dashed). Shown numerical values are for $S_{\rm dS}=76$, while the result of $S_{\rm dS}=760$ is rescaled by Eq. (48) with $c=1/10$, and they overlap very well. Furthermore, the scaling allows to categorize the whole parameter space into three regimes separated by vertical lines. The scrambling timescale turns out to be relevant to the bra-ket wormhole, but the exact location of a theory depends on a true stabilization mechanism. The benchmark parameter in the scrambling regime is marked as a cross.
Our model has four free parameters: $\{S_{\rm dS},\tau_{0},k_{\rm max},m\}$. For Observable 2, $k_{\rm min}$ is also added but is not fixed. In the absence of wormhole stabilization mechanisms, hence without the relation of $S_{\rm dS}$ and $\tau_{0}$, we would first like to explore the possible parameter space and identify interesting ones to focus.
Scaling relation.
Our solution has a scaling relation among the first three of them. The two-point correlations are approximately invariant under the scaling by a constant $c$,
$S_{\rm dS}\,\to\,S_{\rm dS}\cdot c,\quad\tau_{0}\,\to\,\tau_{0}/c,\quad k_{\rm max}^{c}\,\to\,k_{\rm max}^{c}\cdot c.$ (48)
This becomes more exact in the small mass $m\ll 1$ limit. This is numerically demonstrated in Fig. 4, where $k_{\rm max}^{c}$ versus $\tau_{0}$ is shown for two fixed choices $S_{\rm dS}=76$ and $760$, which are related by $c=1/10$. The scaling can be traced back to the $Z_{\rm rat}$ Eq. (45) in the phase transition mechanism, which is approximately calculated at the phase transition as
$\log Z_{\rm rat}\,\propto\,k_{\rm max}^{c}\log\frac{1}{\tau_{0}}\,\simeq\,S_{\rm dS},$ (49)
where the first approximation is numerically demonstrated in the left-bottom panel of Fig. 6; the average value of the mode-function ratio $\sqrt{{\rm Im}{\cal A}_{k}^{HH}/{\rm Im}{\cal A}_{k}^{WH}}$ is proportional to $\log 1/\tau_{0}$ in the small $\tau_{0}\ll 1$ limit. As $1/\tau_{0}$ also sets the scale of $k_{\rm max}^{c}$ (hence, $k_{\rm max}^{c}\propto 1/\tau_{0}$), this proves the approximate scaling relation in Eq. (48).
Scrambling regime.
We can divide the whole parameter space in Fig. 4 into three regimes, separated by vertical lines. The division reflects different timescales of the phase transition, $k_{\rm max}^{c}$. Notably, in the majority of the parameter space, the fast scrambling timescale, $k_{\rm max}^{c}\sim S_{\rm dS}$ (or, $Ht\sim\log S_{\rm dS}$), turns out to be relevant to the bra-ket wormhole physics.
The fast scrambling timescale is when the information thrown into a horizon becomes thermalized enough to be able to leak out of the horizon Hayden:2007cs; Maldacena:2015waa. Why can this timescale be relevant to the two-point correlation across the horizon, via bra-ket wormholes? At first sight, this sounds unreasonable. Wormhole effects are topologically suppressed by $e^{-S_{\rm dS}}$, so that they are expected to be relevant only after a large multiplicity of $e^{S_{\rm dS}}$ is involved Page:1993wv; Arkani-Hamed:2007ryv; in particular, distinguishing a mixed state from a pure state in the Hilbert space of dimension $e^{S}$ requires to measure $e^{S}$ states. This timescale is around the Page time $Ht\sim S_{\rm dS}$. Indeed, in the black hole physics, replica wormholes begin to dominate around this timescale to modify the fine-grained entropy of the black hole Penington:2019kki; Almheiri:2020cfm, completing the Page curve. But correlation functions can be different.
Figure 5: How the Euclidean shortcut via a bra-ket wormhole enhances the antipodal correlation by creating a shorter null geodesic (dotted). For given $\tau_{0}$, the minimum time evolution (modular flow) of $t\simeq\log(4/\tau_{0})$ is needed to create the desired spatial shift $\Delta\theta\simeq\tau_{0}$. This is effectively the same as the shock-wave protocol to create a null geodesic across the horizon Shenker:2013pqa; Gao:2016bin; Maldacena:2017axo; Aalsma:2021kle, hence related to the scrambling physics.
One interpretation of the relevance of the scrambling time involves the shortcut via a Euclidean wormhole shown in the Penrose diagram of Fig. 5. This is a bra-ket wormhole, whose Euclidean region is denoted by the central curly line; we treat our effective wormhole after tracing in this way. When the time evolution (modular flow so that the bra and ket sides are located on the same time values) of two antipodal points is small, no null geodesics can connect them, and the correlation is suppressed. There exists the minimum amount of modular flow $t$ such that the given $\tau_{0}$ can geometrically allow a new null (hence shorter) geodesic between them, enhancing the correlation. At this point, our wormhole metric allows
$Ht\,\simeq\,{\rm arccosh}(1/\sin(\tau_{0}/2))\,\simeq\,\log(4/\tau_{0}),\qquad\Delta\theta\,\simeq\,\tau_{0}.$ (50)
This picture of the Euclidean shortcut via a wormhole is effectively the same as the shock-wave protocol to create a null geodesic across the horizon Shenker:2013pqa; Gao:2016bin; Maldacena:2017axo; Aalsma:2021kle. A signal sent from the South Pole in the past grows chaotically to ${\cal O}(1)$, and finally modifies the geometry to create a shorter null geodesic across the horizon. This growth is exponential near the horizon, yielding the scrambling timescale. Our estimate in Eq. (50) becomes the scrambling time $Ht\sim\log S_{\rm dS}$ if we identify $1/\tau_{0}\sim S_{\rm dS}$. And this is the relation that is observed numerically in the scrambling regime in Fig. 4.
In Ref. Jung:2025rip, authors argued the same result—the scrambling timescale of the phase transition—and confirmed it in random state models. There it was argued based on the competition between the topological suppression $e^{-S}$ and the bra-ket wormhole action $e^{\alpha L}$ for some $\alpha>0$ (due to the independence of the local bra-ket wormhole geometry on the dS boundary length $L$). Interpreting $L\sim e^{Ht}$, they obtained the scrambling time as the phase transition timescale. In our model, a similar competition is responsible for the scrambling timescale. The mode counting in Eq. (45), $Z_{\rm rat}=\prod^{k_{\rm max}}_{k}\sqrt{\frac{{\rm Im}{\cal A}_{k}^{HH}}{{\rm Im}{\cal A}_{k}^{WH}}}\,\sim\,e^{\alpha^{\prime}k_{\rm max}}$ for some $\alpha^{\prime}>0$ (see Eq. (49) and Fig. 6), is actually multiplicative (rather than additive), so that the transition can occur exponentially earlier than the Page time. The reason for the necessary positivity of $\alpha>0$ is the negative Casimir energy in dS+CFT Maldacena:2019cbz; Jung:2025rip, while it is the relaxed KSW or the enhancement of low-$k$ correlations in our solution in dS+QFT.
The fast scrambling timescale is also parametrically close to the timescale estimated to be the lifetime of dS by Trans-Planckian censorship conjecture (TCC) and swampland Bedroya:2019snp, $k_{\rm max}^{c}\sim\sqrt{S_{\rm dS}}$ (or, $Ht\sim\tfrac{1}{2}\log S_{\rm dS}$). The dS phase may cease to continue there or some non-trivial effects may show up to modify this conclusion.
In the rest of the paper, we will focus on the scrambling regime by choosing the following benchmark parameters (marked as a cross in Fig. 4):
$S_{\rm dS}\,=\,76,\quad\tau_{0}\,=\,0.01,\quad m\,=\,0.1,\quad k_{\rm max}\,=\,150.$ (51)
The wormhole stabilization, if available in the future, will only change $\tau_{0}$ for the given $S_{\rm dS}$. The results of the Page regime will be qualitatively similar, except for the possibility of eternal inflation which may require another non-trivial modification. Another reason is to avoid huge computation time for exponentially long path integral for the Page regime.
Lastly, the leftmost region in Fig. 4 is probably invalid. Here, the phase transition occurs too much earlier than the scrambling time. This is probably inconsistent with causality. Thus, we expect that a true wormhole stabilization mechanism will avoid this region.
The scalar mass is less of concern. Since $\phi$ may represent the slow-roll inflaton, this would rather be light. Too heavy ones also spoil the scaling behavior so that the analysis will be complicated. But too light ones, on the other hand, may exhibit remnants of infrared(IR) divergence that arise at $m=0$. Two origins of IR divergences. One is the massless Green’s function, which affects both HH and WH. The other is maximal mixing of $\phi_{1}$ and $\phi_{2}$ in two separate universes ($b_{k}=a_{k}$ in Eq. (37)), which leads to divergence after tracing out one of the universes. This further enhances IR divergence of WH contributions. Thus, we take $m=0.1$ in the benchmark.